OpenStudy (anonymous):

Add. Write your answer as a mixed number in simplest form. 8 3/4 plus 4 5/6

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

yes I can

OpenStudy (anonymous):

To start you need to convert the fraction portions so that they have a common denominator.

OpenStudy (anonymous):

so 4 and 6 have 12 in common

OpenStudy (anonymous):

(8)(3)4+(4)(5)6=24 4+(4)(5) 6=6+ (4)(5) 6=6+206=6+103=283

OpenStudy (anonymous):

28/3

OpenStudy (anonymous):

so 9/12 and 10/12